pooja195 (pooja195):

5x^-2y^10 over 2x^-1(-3x^-3y^-1)^-2 anyone want to help?!

OpenStudy (mathstudent55):

Is this it? \[\frac{ 5x ^{-2 }y ^{10} }{ 2x ^{-1}(-3x ^{-3}y ^{-1})^{-2} }\]

zepdrix (zepdrix):

Oh boy this is a rough one :d Let's start by dealing with all of the negative powers outside of the brackets. Remember what a negative power tells u to do? It tells you to take the reciprocal of the value, FLIP IT. So if something is in the bottom with a negative power, we flip it to the top and change the negative to positive.\[\large \frac{ 5\color{orangered}{x ^{-2 }}y ^{10} }{ 2x ^{-1}(-3x ^{-3}y ^{-1})^{-2} }\]So we'll first apply that idea to this x on top, giving us,\[\large \frac{ 5y ^{10} }{ 2x ^{-1}\cdot \color{orangered}{x^{2}}(-3x ^{-3}y ^{-1})^{-2} }\]See how we FLIPPED IT and it changed to positive?

zepdrix (zepdrix):

Let's do the same thing to this part,\[\large \frac{ 5y ^{10} }{ 2x ^{-1}\cdot x^{2}\color{orangered}{(-3x ^{-3}y ^{-1})^{-2}}}\]Giving us,\[\large \frac{ 5y ^{10}\cdot \color{orangered}{(-3x ^{-3}y ^{-1})^{2}}}{ 2x ^{-1}\cdot x^{2}}\] The -2 changed to a 2 when we brought it up top. Are you following any of this? :o

OpenStudy (mathstudent55):